The width of a straight racetrack is one tenth of the measurement of its length. The
perimeter of the track is 132 metres. What are the dimensions of the length and
width of the track?

Answer by Alan3354(69443)

The width of a straight racetrack is one tenth of the measurement of its length. The perimeter of the track is 132 metres.
What are the dimensions of the length and width of the track?
----------------
L = 10W
L + W = 132/2 = 66
---
10W + W = 66
W = 6
L = 60